# Low-depth quantum symmetrization

### v1

@@ II Preliminaries
Note that if the integer is in $[m]$, then each $a$ or $b$ can be represented with $\log m$ bits. The comparator unitary for $\log m$ qubits can be implemented in $O(\log\log m)$ depth, as shown in Appendix B.2 of Ref. Berry et al. (2018). When analyzing the circuit depth, we ignore this double-logarithmic overhead for simplicity.
@@
Multiple sorting networks with sub-linear depth have been proposed, including the well-known bitonic sort Batcher (1968) and merge sort Stone (1971), both featuring $O(\log^{2}n)$ depth for a list of $n$ elements (and $N_{a}=O(n\log^{2}n)$ ancilla qubits when quantized). As far as we are aware, the only known $O(\log n)$-depth sorting network is the AKS algorithm proposed in Ref. Ajtai et al. (1983). However, this algorithm is impractical due to its large constant factor.
@@
In this paper, we employ the AKS algorithm to design our symmetrization algorithms, but we emphasize that all sorting networks we use can be replaced by bitonic or merge sort if the number of elements to be sorted is not too large. While the asymptotic depth using bitonic or merge sort will have another factor of $\log n$, the performance may be better in practice. Since all four operations ($\mathrm{SORT},\mathrm{UNSORT},\mathrm{SHUFFLE},\mathrm{UNSHUFFLE}$) only use as many gates as running a quantized sorting network, they can all be performed in $O(\log n)$ depth.
II.2 SIL symmetrization algorithms
Symmetrization and antisymmetrization algorithms are fundamental building blocks of first-quantized quantum simulation for bosonic and fermionic systems. Abrams and Lloyd (1997) proposed the first quantum (anti-)symmetrization algorithm. This procedure has $\operatorname{poly}(n)$ depth. Berry et al. (2018) improved their algorithm by introducing $\log(n)$-depth sorting networks and a $\log(n)$-depth method for generating random permutations. However, both algorithms are only applicable to SIL inputs, although this was not explicitly pointed out in their papers. In this subsection, we briefly introduce the algorithm of Berry et al. (2018), as it is a building block of our NSIL symmetrization algorithm.
@@ II.2 SIL symmetrization algorithms
where $\ket{\perp}$ is the superposition of repetitive $r$s. Note that Berry et al. (2018) include a measure-and-postselect step to remove the repetitive part. We can also choose to use $a\geq 3$ and keep the repetitive part since it only leads to a negligible error.
Next, sort register 2 using the $O(\log n)$-depth sorting network and store the record in register 3:
@@
In fact, the last step of the Abrams and Lloyd (1997) algorithm is also unsorting the input list using the superposition of all records in $S_{n}$, but the $\mathrm{UNSORT}$ operation is implemented based on $\operatorname{poly}(n)$-depth sorting algorithms.
@@
This algorithm, first proposed in Hillis and Steele Jr. (1986), is based on the divide-and-conquer paradigm and uses $O(\log^{2}n)$ circuit depth.
@@
The subroutine $D(h,t)$ can be implemented in $O(\log n)$ depth since it involves adding $d_{(h+t)/2}$ to $(t-h)/2$ registers. Hence the total circuit depth is $O(\log^{2}n)$.
@@ III Symmetrization of non-strictly increasing lists
In this section, we consider symmetrizing non-strictly increasing that may contain duplicated items, i.e., $\mathrm{NSIL}$s. In Section III.1, we introduce some useful subsets and subgroups of $S_{n}$ related to $\mathrm{NSIL}$s. In Section III.2, we describe the behaviors of $\mathrm{SORT}$ and $\mathrm{UNSORT}$ on NSILs, and explain why the SIL symmetrization algorithms do not apply to $\mathrm{NSIL}$s. In Section III.3, we describe our new algorithm for symmetrizing |$\ket{l}$ where $l\in\mathrm{NSIL}$, and finally in Section III.4, we extend it to the case of superposed inputs, i.e., $\sum_{l}\alpha_{l}\ket{l}$.
@@ III.1 Degenerate permutations for NSILs
There is a kind of degeneracy of permutations for $l\in[m]^{n}$ with repetitions (in this paper, we mostly study $l\in\mathrm{NSIL}_{n}^{m}\subset[m]^{n}$): different permutations $\sigma_{1},\sigma_{2}\in S_{n}$ may lead to the same permuted list, i.e., $\sigma_{1}(l)=\sigma_{2}(l)$. In this subsection, we show that this degeneracy is the fundamental reason why the Berry et al. (2018) algorithm does not apply to NSILs. In particular, let $H_{l}$ denote a subset of the permutation group $S_{n}$ which contains only permutations preserving the input $l\in[m]^{n}$, i.e.,
@@ III.2 Quantum sorting networks for NSILs
Recall that when introducing sorting-network-related operations in Section II, $\mathrm{SORT}$ is designed only to sort $\sigma(l)$ where $l$ is strictly increasing, and $\mathrm{UNSORT}$ is also only designed to act on SILs. Here we discuss what happens if both operations are applied to NSILs. When $l$ contains duplicates, the $\mathrm{SORT}$ operation still transforms $\sigma(l)$ to $l$, but what is stored in the record depends on the stability of the sorting algorithm. Recall from Section II that we are free to choose the sorting algorithm. A sorting algorithm is called stable if it preserves the relative order of duplicated elements. Therefore, sorting any permutation in a given coset using a stable algorithm produces the record corresponding to the canonical element we picked ($R_{l}$), i.e., the record of sorting $\sigma(l)$ will be $\operatorname{rec}(\sigma^{\prime})$ where $\sigma^{\prime}(l)=\sigma(l)$ and $\sigma^{\prime}\in R_{l}$. (Recall that $\operatorname{rec}(\sigma)$ is the record obtained from sorting $\sigma(l_{\mathrm{S}})$ where $l_{\mathrm{S}}$ can be any $\mathrm{SIL}$.) On the other hand, unstable sorting networks shuffle the relative order of duplicated elements, so the record of sorting $\sigma(l)$ will be $\operatorname{rec}(\sigma^{\prime}\circ h^{\prime})$ where $h^{\prime}$ is an unknown permutation in $H_{l}$ depending on $\sigma^{\prime}$. In this paper, we consider general unstable sorting networks, since many efficient sorting networks, including bitonic sort Ajtai et al. (1983), are unstable.
@@ III.3 NSIL symmetrization for a single input
In this subsection, we describe our algorithm for symmetrizing NSILs for input state |$\ket{l}$ using our observations mentioned in Sections III.1 and III.2. The algorithm extensively uses all four sorting network operations introduced in Section II. We use three registers (the latter two are ancilla registers): register 1 stores the input NSIL $l$, register 2 stores all records for the sorting algorithm and is initialized as all zero, and register 3 stores the superposition of all $\sigma(1,2,\dots,n)$ initially, which can be generated using either the Berry et al. (2018) SIL symmetrization algorithm (with $O(\log n)$ depth) or our new algorithm described in Appendix A (with $O(\log^{3}(n))$ depth). Register 3 is used as a platform to mediate multiplications of permutations, which eventually allows us to extract the subgroup structure corresponding to NSILs.
As the first step, we sort each $\sigma(1,2,\dots,n)$ reversibly using sorting networks and store the records of swaps:
@@
We notice that the strings in register 3 are now only superpositions of $h(12\dots n)$ over $h\in H_{l}$, and are independent of which permutation is applied in register 1. Therefore, if the input is a single NSIL basis state |$\ket{l}$, one can simply discard registers 2 and 3, and the state in register 1 is $\sum_{\sigma\in S_{n}}\ket{\sigma(l)}$. The output fidelity of single-input symmetrization is $1$ if the resource state $\sum_{\sigma}\ket{\sigma(12\dots n)}$ is prepared exactly. The Berry et al. (2018) algorithm can produce this state with high probability.
III.4 NSIL symmetrization for a superposed input
@@
from which we cannot obtain |$\ket{\psi_{\mathrm{out}}}$ by simply discarding register $3$. Hence, we must erase the scratch qubits in register 3 for all possible input strings $l$ coherently. This can be done by inverting the circuit preparing the superposition of the $H_{l}$ subgroup ($\sum_{h\in H_{l}}\ket{h(12\dots n}$) from |$\ket{l}$. Here we describe how to prepare this subgroup equal superposition by adapting the SIL symmetrization algorithm of Berry et al. (2018). Note we can also use our new SIL algorithm for this task, as shown in Appendix A.
Since $h\in H_{l}$ contains only permutations between repeated elements, non-repeated elements should stay where they are. We can modify the Berry et al. (2018) algorithm to only generate this kind of permutations. When preparing the initial state for register 2, we simply need to change the lower and upper limits of the $r_{i}$s such that
|$\ket{\phi_{2}}\propto\bigotimes_{i=1}^{n}\left(\sum_{r_{i}=(l_{i}-1)f(n)}^{l_{i}f(n)-1}\ket{r_{i}}\right).$ (43)
To see why this state corresponds to only permutations in $H$, we observe that if $l_{i}<l_{j}$, then $r_{i}<r_{j}$, hence no permutation happens between the $i$th item and the $j$th item. If $l_{i}=l_{j}$, then $r_{i}$ may be less than, equal to, or greater than $r_{j}$, hence permutations are allowed between $i$ and $j$.
Since the greatest possible value of $r_{i}$ is now $f(n)\max(l_{i})$ and $\max(l_{i})\leq m$, the number of qubits used is $n\log(mn^{a})=O(n\log(m)+an\log(n))$.
Finally, we analyze the fidelity of the symmetrized states produced by this algorithm acting on superposed inputs. Observe that the probability of obtaining repetitive $r$ from measuring |$\ket{\phi_{2}}$ in Equation (43) depends on the pattern of repetition in $l$. For two extreme examples, if $l\in\mathrm{SIL}$, then there will be no repetitive $r$ in |$\ket{\phi_{2}}$; however, if $l=\ket{00\dots 0}$ or any integer repeated $n$ times, then the probability of obtaining a repetitive $r$ is $O(n^{-a+2})$, just as in the original SIL symmetrization algorithm. As a result, if the input state is some superposition $\sum_{l\in\mathrm{NSIL}}\alpha_{l}\ket{l}$ of NSILs with different repetition patterns, then the output state will not be exactly $\sum_{l\in\mathrm{NSIL}}\alpha_{l}\ket{\psi_{\mathrm{out}}(l)}$, but will have some slightly different amplitudes, $\sum_{l\in\mathrm{NSIL}}\alpha^{\prime}_{l}\ket{\psi_{\mathrm{out}}(l)}$. Since $\alpha^{\prime}_{l}$ arises from measuring |$\ket{\phi_{2}}$ and getting no repetition in Equation (43), $\alpha^{\prime}/\alpha$ is the square root of the probability of having repetition in that step rescaled for normalization, which is a real number. We can therefore derive the following lower bound:
$\left|\frac{\alpha^{\prime}_{l}}{\alpha_{l}}\right|^{2}\geq\frac{\left[1-O(n^{-a+2})\right]|\alpha_{l}|^{2}}{|\alpha_{l}|^{2}}=1-O(n^{-a+2}).$ (44)
Note that even the extreme case $l=\ket{00\dots 0}$ usually cannot saturate this lower bound, since $\sum_{l\in\mathrm{NSIL}}\alpha^{\prime}_{l}\ket{\psi_{\mathrm{out}}(l)}$ is also a normalized state.
We can now lower bound the output fidelity using $\sum_{l}|\alpha_{l}|^{2}=1$:
$\left|\sum_{l\in\mathrm{NSIL}}\alpha^{*}_{l}\alpha^{\prime}_{l}\right|^{2}=\left(\sum_{l\in\mathrm{NSIL}}\alpha^{*}_{l}\alpha^{\prime}_{l}\right)^{2}\geq 1-O(n^{-a+2}).$ (45)
Therefore, the output fidelity of the $O(\log n)$-depth NSIL symmetrization algorithm is $1-O(n^{-a+2})$ for superposed inputs. This error is non-zero but negligible.
IV Efficient conversion from second quantization to first quantization
To convert any given second-quantized state to its first-quantized counterpart, we first transform the second-quantized state to an NSIL before employing the (anti-)symmetrization algorithm on the increasing lists. To do so, we design a novel reversible circuit that takes as input a list of occupation numbers (i.e., a second-quantized representation) and outputs an increasing list of positions (or modes) occupied, where each mode is assigned with a number written in its binary representation. More specifically, our circuit transforms $W_{0}:=n_{0}n_{1}\dots n_{m-1}$ (where $n_{i}\geq 0$ is the occupation number of mode $i$) into $W_{f}:=0^{n_{0}}1^{n_{1}}\dots(m-1)^{n_{m-1}}$. Our approach is based on reversible sorting networks and takes $O(\log^{2}n)$ depth where $n=\sum_{i=0}^{m-1}n_{i}$. The main idea for converting $W_{0}$ to $W_{f}$ reversibly is to first generate $n$ empty registers; then use sorting networks to move $n_{i}$ of them in between of the $(i-1)$th register and the $i$th register, for all $i\in[m]$; and use the new positions of the empty registers to compute the index $i$ locally for each of them. In more detail, the algorithm proceeds as follows.
$\tilde{W}_{1}$ $=(0,n_{0},n_{0})(1,n_{1},n_{0}+n_{1})(2,n_{2},n_{0}+n_{1}+n_{2})$ (47)
$\quad\dots(m-1,n_{m-1},n).$
$\tilde{W}_{2}$ $=(0,n_{0},n_{0})(1,n_{1},n_{0}+n_{1})\dots$ (48)
$\quad(m-1,n_{m-1},n)(m,0,0)(m,0,1)\dots(m,0,n-1).$
  • $a_{1}<a_{2}$;
  • $a_{1}=a_{2}=m$ AND $c_{1}<c_{2}$.
Now, using reversible sorting networks, we sort $\tilde{W}_{2}$ to obtain
$\quad\dots(m,0,n_{0}+n_{1}-1)(1,n_{1},n_{0}+n_{1})(m,0,n_{0}+n_{1})$
$\quad\dots(m,0,n-1)(m-1,n_{m-1},n),$
  • $c_{1}<c_{2}$,
  • $c_{1}=c_{2}$ AND $a_{1}<a_{2}$.
Now, we have allocated $n_{i}$ such elements (with first element being $m$) between $(i-1,n_{i-1},\sum_{j=0}^{i-1}n_{j})$ and $(i,n_{i},\sum_{j=0}^{i}n_{j})$, and we need to assign the value $i$ to each of them. For the $k$th element in $\tilde{W}_{3}$, if the first value is $m$ (meaning that this is a newly added element) and the third value is $i_{k}$ (the index of new elements), then the corresponding $i$ value is simply $k-i_{k}$ (the difference between the index in $\tilde{W}$ and the index in all new elements). We add this value to the second value for each new element, obtaining
$\tilde{W}_{4}$ $=(m,0,0)(m,0,1)\dots(m,0,n_{0}-1)(0,n_{0},n_{0})$ (50)
$\quad(m,1,n_{0})\dots(m,1,n_{0}+n_{1}-1)(1,n_{1},n_{0}+n_{1})$
$\quad(m,2,n_{0}+n_{1})\dots(m,m-1,n-1)$
$\quad(m-1,n_{m-1},n).$
$\tilde{W}_{5}$ $=(0,n_{0},n_{0})(1,n_{1},n_{0}+n_{1})$ (51)
$\quad\dots(m-1,n_{m-1},n)(m,0,0)\dots.$
$\tilde{W}_{7}$ $=(m,0,0)\dots(m,0,n_{0}-1)(0,0,n_{0})(m,1,n_{0})\dots$ (53)
$\quad(m,1,n_{0}+n_{1}-1)(1,0,n_{0}+n_{1})(m,2,n_{0}+n_{1})$
$\quad\dots(m,m-1,n-1)(m-1,0,n).$
Next, we notice that $(i,0,\sum_{j=0}^{i}n_{i})$ is always right after $(m,i,\sum_{j=0}^{i}n_{i}-1)$ for all $i$. We can therefore subtract all $\sum_{j=0}^{i}n_{i}$ values and obtain
$\tilde{W}_{8}$ $=(m,0,0)\dots(m,0,n_{0}-1)(0,0,0)(m,1,n_{0})\dots$ (54)
$\quad(m,1,n_{0}+n_{1}-1)(1,0,0)(m,2,n_{0}+n_{1})\dots$
$\quad(m,m-1,n-1)(m-1,0,0).$
Next, we can sort $\tilde{W}_{8}$ again under $\Gamma_{W_{2}}$ to obtain
$\tilde{W}_{f}$ $=(0,0,0)\dots(m-1,0,0)(m,0,0)\dots$ (55)
$\quad(m,0,n_{0}-1)(m,1,n_{0})\dots(m,1,n_{0}+n_{1}-1)\dots$
$\quad(m,m-1,n-1).$
$W_{f}=0^{n_{0}}1^{n_{1}}\dots(m-1)^{n_{m-1}}.$ (56)
Since all operations in this process are either reversible sorting or reversible local data modification, the whole process takes $\operatorname{poly}(\log n)$ depth and can be implemented as a reversible circuit.
V Application: quantum inteferometric imaging system
V.1 QFT-based single-photon interferometry
We review quantum telescope arrays Gottesman et al. (2012); Khabiboulline et al. (2019a) for the low-photon-rate scenario with the help of the schematic diagram in Fig. 4. For simplicity, we work in an idealized model of a 2-dimensional world where all $m$ telescopes are equally spaced in a line. Each telescope $T_{0},T_{1},\dots,T_{m-1}$ has a lens (or mirror) to collect photons and a well-isolated photon detector equipped with a bosonic quantum memory (since photons are bosons). The distance between adjacent telescopes is $d$. The photon source (which may be a star, galaxy, etc.) is denoted by $S$ and is assumed to be infinitely far away from the telescope array, so its position can be described using a single parameter $\theta$, the angle between $T_{0}S$ and $SP$ where $P$ is the nearest point to $S$ on the telescope plane. Since both $d$ and $m$ are finite, the following angle relation can be derived from trigonometry:
@@ V.1 QFT-based single-photon interferometry
$a^{\dagger}(\theta)\ket{00\dots 0}$ (59)
[…]
We can now perform the QFT on this state to obtain
Observe that if
$\frac{d\sin\theta}{\lambda}-\frac{k}{m}=v$ (63)
$k\left(\frac{\lambda}{md}\right)=\sin\theta-\frac{v\lambda}{d}.$ (64)
Note that in the discussions of quantum telescope arrays in this paper, we assume there are no atmospheric fluctuations that may produce local phase differences. This may be reasonable if the telescope array is space-based, but for ground-based telescopes, such phases are inevitable and may affect image quality, since the phase difference between the $i$th and the $j$th telescopes is no longer strictly $2\pi d(j-i)\sin\theta/\lambda$. These fluctuations can potentially be corrected using classical methods such as adaptive optics Beckers (1993), as mentioned in Ref. Khabiboulline et al. (2019b).
V.2 Multiple-photon interferometry
  1. 1.
    the number of detectors $m$ is large;
  2. 2.
    the area of each aperture is large;
  3. 3.
    the field of view is large (which can be achieved when $d$ is very small);
  4. 4.
    the photon wavelength is long; or
  5. 5.
    the source is bright.
Now, we outline how our NSIL symmetrization algorithm makes it possible to extract all $\theta_{j}$ from |$\ket{\psi(\theta_{1},\dots,\theta_{n})}$ using only one copy of the state with perfect success probability when using Result 2, or with high confidence when using Result 1. Let
$\phi_{i,j}:=\frac{2\pi dj\sin\theta_{i}}{\lambda}.$ (66)
The explicit multiple-photon state is
$\prod_{j=1}^{n}a^{\dagger}(\theta_{j})\ket{00\dots 0}$ (67)
As with the single-photon case, we convert to the first-quantized representation in order to apply the quantum Fourier transform. To do so, as described in Section IV, we first transform |$\ket{n_{0}n_{1}\dots n_{m-1}}$ into a list $l$ of telescope indices in increasing order. This transformation circuit has $O(\log^{2}n)$ depth and uses $O(n\log m)$ ancilla qubits. As a simple example, if $m=3$, $n=4$, and $r=2101$, then the state in Equation (68) is |$\ket{121}$, which will be transformed to |$\ket{l(r)}=\ket{0112}$. Here $l(r)=0112$ is an NSIL as it is the sorted version of $r$. We also notice that the factor $\sqrt{\prod_{i=0}^{m-1}n_{i}!}$ is just $\sqrt{|H_{l(r)}|}$, square root of the order of the subgroup of $S_{n}$ preserving $l(r)$. We can thus write the transformed unnormalized state as
Next, we can apply the NSIL symmetrization algorithm to obtain the unnormalized first-quantized state
|$\ket{\psi_{\mathrm{S}}(\theta_{1},\dots,\theta_{n})}$ (70)
Finally, we rewrite
|$\ket{\psi_{\mathrm{S}}(\theta_{1},\dots,\theta_{n})}$ (74)
$\mathrm{QFT}^{n}\ket{\psi_{\mathrm{S}}(\theta_{1},\dots,\theta_{n})}=\sum_{\sigma\in S_{n}}\bigotimes_{i=1}^{n}\ket{\theta_{\sigma(12\dots n)_{i}}}.$ (75)
It is also worth mentioning that instead of directly producing images, quantum telescope arrays can offer better accuracy in estimating specific parameters, such as the distance between binary stars, as introduced in Ref. Sajjad et al. (2024). Our multiple-photon interferometric imaging system can potentially extend the list of high-precision parameter estimations.
VI Summary & Discussion
In this paper, we studied the general quantum symmetrization problem. We identified subtleties with repetitions in the input list and pointed out issues with previous SIL symmetrization algorithms when acting on NSILs. We then proposed the first quantum algorithms for symmetrizing NSILs using sorting networks. These algorithms leverage the coset decomposition of the symmetric group, use an ancilla register to mediate multiplications of permutations elements, and adapt SIL symmetrization algorithms as subroutines. The depth-optimal algorithm features $O(\log n)$ depth and uses $O(n\log m)$ ancilla qubits for NSILs of $n$ integers with value up to $m$. We also proposed an $O(\log^{2}n)$-depth to transform between second-quantized states and $\mathrm{NSIL}$s, which allows for efficient conversion between second-quantized states and first-quantized states.
We distinguish between the scenario in our paper and in other algorithmic applications of symmetrization. It is well known that the graph isomorphism problem can be solved by creating a superposition over all permutations of a given graph Ambainis et al. (2011). However, our results cannot be used in that context since there is no increasing input guarantee in the setting of graph isomorphism. Indeed, if the vertices of a graph could be ordered in a canonical way, this would solve graph isomorphism.
Our algorithm requires all-to-all connectivity due to the sorting networks used, just as in Ref. Berry et al. (2018). A natural open problem is to find low-depth sorting networks with restricted connectivity, which could make our algorithm much more favorable for near-term devices. It might also be possible to reduce the circuit depth of the exact symmetrization algorithm in Result 2 or increase the fidelity of the $O(\log n)$-depth algorithm in Result 1.
Acknowledgments
Appendices
In the appendices, we describe the exact NSIL symmetrization algorithm mentioned in Result 2. In Appendix A, we describe the exact SIL symmetrization algorithm based on quantization of the algorithm of Alonso and Schott (1996). In Appendix B, we describe the exact NSIL symmetrization algorithm.
Appendix A An exact SIL symmetrization algorithm

### v2

@@ II Preliminaries
Note that if the integer is in $[m]$, then each $a$ or $b$ can be represented with $\log m$ bits. The comparator unitary for $\log m$ qubits can be implemented in $O(\log\log m)$ depth, as shown in Appendix B.2 of Ref. [4]. When analyzing the circuit depth, we ignore this double-logarithmic overhead for simplicity.
@@
Multiple sorting networks with sub-linear depth have been proposed, including the well-known bitonic sort [29] and merge sort [30], both featuring $O(\log^{2}n)$ depth for a list of $n$ elements (and $N_{a}=O(n\log^{2}n)$ ancilla qubits when quantized). As far as we are aware, the only known $O(\log n)$-depth sorting network is the AKS algorithm proposed in Ref. [31]. However, this algorithm is impractical due to its large constant factor.
@@
In this paper, we employ the AKS algorithm to design our symmetrization algorithms, but we emphasize that all sorting networks we use can be replaced by bitonic or merge sort if the number of elements to be sorted is not too large. While the asymptotic depth using bitonic or merge sort will have another factor of $\log n$, the performance may be better in practice. Since all four operations ($\mathrm{SORT},\mathrm{UNSORT},\mathrm{SHUFFLE},\mathrm{UNSHUFFLE}$) only use as many gates as running a quantized sorting network, they can all be performed in $\tilde{O}(\log n)$ depth.
II.2 Berry et al. SIL symmetrization algorithms
Symmetrization and antisymmetrization algorithms are fundamental building blocks of first-quantized quantum simulation for bosonic and fermionic systems. Abrams and Lloyd [2] proposed the first quantum (anti-)symmetrization algorithm. This procedure has $\operatorname{poly}(n)$ depth. Berry et al. [4] improved their algorithm by introducing $\log(n)$-depth sorting networks and a $\log(n)$-depth method for generating random permutations. However, both algorithms are only applicable to SIL inputs, although this was not explicitly pointed out in their papers. In this subsection, we briefly introduce the algorithm of Berry et al. [4], as it is a building block of our NSIL symmetrization algorithm.
@@
where $\ket{\perp}$ is the superposition of repetitive $r$s. Note that Berry et al. [4] include a measure-and-postselect step to remove the repetitive part. We can also choose to use $a\geq 3$ and keep the repetitive part since it only leads to a negligible error.
Next, sort register 2 using the $\tilde{O}(\log n)$-depth sorting network and store the record in register 3:
@@
In fact, the last step of the Abrams and Lloyd [2] algorithm is also unsorting the input list using the superposition of all records in $S_{n}$, but the $\mathrm{UNSORT}$ operation is implemented based on $\operatorname{poly}(n)$-depth sorting algorithms.
@@
This algorithm, first proposed in [32], is based on the divide-and-conquer paradigm and uses $O(\log^{2}n)$ circuit depth.
@@
The subroutine $D(h,t)$ can be implemented in $O(\log n)$ depth since it involves adding $d_{(h+t)/2}$ to $(t-h)/2$ registers. Hence the total circuit depth is $\tilde{O}(\log^{2}n)$.
@@ III Symmetrization of non-strictly increasing lists
In this section, we consider symmetrizing non-strictly increasing that may contain duplicated items, i.e., $\mathrm{NSIL}$s. In Section III.1, we introduce some useful subsets and subgroups of $S_{n}$ related to $\mathrm{NSIL}$s. In Section III.2, we describe the behaviors of $\mathrm{SORT}$ and $\mathrm{UNSORT}$ on NSILs, and explain why the SIL symmetrization algorithms do not apply to $\mathrm{NSIL}$s. In Section III.3, we describe our new algorithm for symmetrizing a single $l$ where $l\in\mathrm{NSIL}$. We defer the algorithm for superpositions of $\mathrm{NSIL}$s (i.e., inputs of the form $\sum_{l}\alpha_{l}\ket{l}$) to Section IV.
@@ III.1 Degenerate permutations for NSILs
There is a kind of degeneracy of permutations for $l\in[m]^{n}$ with repetitions (in this paper, we mostly study $l\in\mathrm{NSIL}_{n}^{m}\subset[m]^{n}$): different permutations $\sigma_{1},\sigma_{2}\in S_{n}$ may lead to the same permuted list, i.e., $\sigma_{1}(l)=\sigma_{2}(l)$. In this subsection, we show that this degeneracy is the fundamental reason why the Berry et al. [4] algorithm does not apply to NSILs. In particular, let $H_{l}$ denote a subset of the permutation group $S_{n}$ which contains only permutations preserving the input $l\in[m]^{n}$, i.e.,
@@ III.2 Quantum sorting networks for NSILs
Recall that when introducing sorting-network-related operations in Section II, $\mathrm{SORT}$ is designed only to sort $\sigma(l)$ where $l$ is strictly increasing, and $\mathrm{UNSORT}$ is also only designed to act on SILs. Here we discuss what happens if both operations are applied to NSILs. When $l$ contains duplicates, the $\mathrm{SORT}$ operation still transforms $\sigma(l)$ to $l$, but what is stored in the record depends on the stability of the sorting algorithm. Recall from Section II that we are free to choose the sorting algorithm. A sorting algorithm is called stable if it preserves the relative order of duplicated elements. Therefore, sorting any permutation in a given coset using a stable algorithm produces the record corresponding to the canonical element we picked ($R_{l}$), i.e., the record of sorting $\sigma(l)$ will be $\operatorname{rec}(\sigma^{\prime})$ where $\sigma^{\prime}(l)=\sigma(l)$ and $\sigma^{\prime}\in R_{l}$. (Recall that $\operatorname{rec}(\sigma)$ is the record obtained from sorting $\sigma(l_{\mathrm{S}})$ where $l_{\mathrm{S}}$ can be any $\mathrm{SIL}$.) On the other hand, unstable sorting networks shuffle the relative order of duplicated elements, so the record of sorting $\sigma(l)$ will be $\operatorname{rec}(\sigma^{\prime}\circ h^{\prime})$ where $h^{\prime}$ is an unknown permutation in $H_{l}$ depending on $\sigma^{\prime}$. In this paper, we consider general unstable sorting networks, since many efficient sorting networks, including bitonic sort [31], are unstable.
@@ III.3 NSIL symmetrization for a single input
In this subsection, we describe our algorithm for symmetrizing NSILs for input state |$\ket{l}$ using our observations mentioned in Sections III.1 and III.2. The algorithm extensively uses all four sorting network operations introduced in Section II. We use three registers (the latter two are ancilla registers): register 1 stores the input NSIL $l$, register 2 stores all records for the sorting algorithm and is initialized as all zero, and register 3 stores the superposition of all $\sigma(12\dots n)$ initially, which can be generated using either the Berry et al. [4] SIL symmetrization algorithm (with $\tilde{O}(\log n)$ depth) or our new algorithm described in Section IV.1 (with $\tilde{O}(\log^{3}(n))$ depth). Register 3 is used as a platform to mediate multiplications of permutations, which eventually allows us to extract the subgroup structure corresponding to NSILs.
As the first step, we sort each $\sigma(12\dots n)$ reversibly using sorting networks and store the records of swaps:
@@
We notice that the strings in register 3 are now only superpositions of $h(12\dots n)$ over $h\in H_{l}$, and are independent of which permutation is applied in register 1. Therefore, if the input is a single NSIL basis state |$\ket{l}$, one can simply discard registers 2 and 3, and the state in register 1 is $\sum_{\sigma\in S_{n}}\ket{\sigma(l)}$. The output fidelity of single-input symmetrization is $1$ if the resource state $\sum_{\sigma}\ket{\sigma(12\dots n)}$ is prepared exactly. The Berry et al. [4] algorithm can produce this state with high probability, as mentioned in Section II.2.
IV NSIL symmetrization for a superposed input
@@
from which we cannot obtain |$\ket{\psi_{\mathrm{out}}}$ by simply discarding register $3$. Hence, we must erase the scratch qubits in register 3 for all possible input strings $l$ coherently. This can be done by inverting the circuit preparing the superposition of the $H_{l}$ subgroup ($\sum_{h\in H_{l}}\ket{h(12\dots n}$) from |$\ket{l}$. Therefore, in this section, we design a unitary to implement the following subgroup superposition preparation procedure:
@@
In this section, we first design a $\operatorname{poly}(\log n)$-depth quantum algorithm for SIL symmetrization with perfect success probability in Section IV.1, which is used as a subroutine to solve the subgroup superposition preparation problem in Section IV.2.
IV.1 An exact SIL symmetrization algorithm